what if we assume son is as old as his father is today then there ages total will be 126 and if father is as old as his son is today then there ages total will be 38. find there present ages
Answer by ankor@dixie-net.com(22740)

what if we assume son is as old as his father is today then there ages total will be 126
s + 2f = 126
if s = f, then
f + 2f = 126
f = 126/3
f = 42 is father's present age
:
and if father is as old as his son is today then there ages total will be 38. find there present ages
2s = 38
s = 19 is son's present age